# X-IONet: Cross-Platform Inertial Odometry Network with Dual-Stage Attention

### v1

@@ III-C2 State Propagation
$\hat{b}_{g_{(n+1)}}=\hat{b}_{g_{n}}+\sigma_{g_{n}}$ (12)
$\hat{b}_{a_{(n+1)}}=\hat{b}_{a_{n}}+\sigma_{a_{n}}$ (13)
Where $\exp_{SO3}$ is the exponential map for the rotation group, and $\Delta t$ is the sampling period of the inertial data. $\boldsymbol{g}$ represents the gravity acceleration in the world frame. $\sigma$ donates Gaussian-distributed noise. The propagation of the error state is represented by a linearized matrix:
$\boldsymbol{\hat{x}}_{n+1}=\boldsymbol{A}_{n}^{x}\boldsymbol{\tilde{x}}_{n}+\boldsymbol{B}_{n}^{x}\boldsymbol{n}_{n}$ (14)
Where $\boldsymbol{n}_{k}$ is a vector containing all noise, and $\boldsymbol{A}_{k}^{x}$ , $\boldsymbol{B}_{k}^{x}$ are the state transition matrix and noise input matrix, respectively. The covariance propagation model is:
$\boldsymbol{P_{n+1}}=\boldsymbol{A}_{n}\boldsymbol{P}_{n}\boldsymbol{A}_{n}^{T}+\boldsymbol{B}_{n}\boldsymbol{W}_{n}\boldsymbol{B}_{n}^{T}$ (15)
where $P_{n}$ and $W_{n}$ denote the covariance matrices of the system state and the noise at time $n$, respectively.
@@ III-C3 State Update
$h(\boldsymbol{X})=\mathbf{R}^{T}(\mathbf{{}^{w}}\boldsymbol{p}_{j}-\mathbf{{}^{w}}\boldsymbol{p}_{i})=\boldsymbol{\hat{d}}_{ij}+\boldsymbol{\sigma}_{ij}$ (16)
where $\mathbf{R}^{T}$ represents the transformation that converts the displacement difference in the world frame to the local frame. The corresponding Jacobian matrix is given by:
$\mathbf{H}_{\theta_{i}}=\frac{\partial{h(\boldsymbol{X})}}{\partial{\tilde{\theta}_{i}}}=\mathbf{\hat{R}}^{T}\lfloor\mathbf{{}^{w}}\boldsymbol{p}_{j}-\mathbf{{}^{w}}\boldsymbol{p}_{i}\rfloor_{\times}$ (17)
$\mathbf{H}_{\delta\tilde{\mathbf{p}}}=\frac{\partial{h(\boldsymbol{X})}}{\partial{\delta\tilde{p}_{j}}}=\mathbf{\hat{R}}^{T}$ (18)
Using the measurement Jacobian matrix $\mathbf{H}$ and the covariance $\boldsymbol{\Sigma}$ output by the neural network, the Kalman gain $\mathbf{K}$ is calculated to update the state $\mathbf{S}$ and covariance:
@@
$\mathbf{P}\leftarrow(\boldsymbol{I}-\mathbf{K}\mathbf{H})\boldsymbol{P}(\boldsymbol{I}-\mathbf{K}\mathbf{H})^{T}+\mathbf{K}\hat{\boldsymbol{\Sigma}}_{ij}\mathbf{K}^{T}$ (21)
where $\oplus$ denotes the addition operation on the $SO(3)$ group when updating pose state variables, while other state variables use ordinary addition.
@@ IV EXPERIMENT
TABLE II: Results comparison of different models on RoNIN dataset.
Model RoNIN_seen RoNIN_unseen
ATE RTE ATE RTE
RoNIN-ResNet 4.11 3.28 5.26 4.62
RoNIN-TCN 4.31 5.02 6.74 5.10
RONIN-LSTM 6.19 4.56 7.61 5.87
TLIO 4.74 3.96 6.11 4.75
IMUNet 6.56 5.90 8.35 6.02
MobileNet 6.89 5.09 8.80 6.96
MnasNet 3.96 2.80 6.08 5.03
EfficientNet 4.59 3.46 6.25 4.84
CTIN 4.34 3.08 5.25 4.54
Ours 3.10 2.42 4.79 4.05
@@ IV-A1 Datasets
We evaluated the proposed method on both pedestrian and quadruped robot inertial datasets. For pedestrians, we adopted RoNIN [14], a large-scale, high-quality dataset containing over 42.7 hours of IMU data with ground-truth trajectories. We followed the official training, validation, and test splits. The test set is divided into Seen and Unseen categories, with 32 trajectories each, corresponding to subjects included in or excluded from the training set, respectively.
For quadruped robots, we collected a dataset using the Unitree Go2 EDU platform, as shown in Fig. 3. Ground-truth trajectories and the rotation $^{w}_{b}\mathbf{R}\in SO(3)$ between the IMU frame and the world frame were obtained via LiDAR odometry from an onboard Unitree L1 LiDAR. Since the robot’s IMU orientation is unknown relative to the body frame, we mounted an iPhone 14 Pro Max on top of the robot and performed attitude calibration. We obtained the rotation relationship $^{b}_{i}\mathbf{R}\in SO(3)$ between the IMU frame and the robot body frame. The phone’s IMU provided inertial measurements at 100 Hz, which were resampled to 200 Hz along with the ground-truth trajectories using linear interpolation. We recorded three difficulty levels—Easy, Medium, and Hard—each containing 10 sequences with a total duration of 1.5 hours. The data were split into training, validation, and test sets at a ratio of 3:1:1.
Figure 3: The quadruped robot used in the experiments.
Figure 4: Trajectory comparisons of partial experimental results. The top three trajectory plots illustrate the comparisons of different methods on the Go2_Easy_01, Go2_Medium_02, and Go2_Hard_02, while the bottom three plots correspond to the RoNIN_a037_1, RoNIN_a038_2, and RoNIN_a044_2. In the figures, darker and thicker trajectories indicate a closer match to the ground-truth motion paths. Each image highlights selected regions with magnified details to better illustrate the differences among the methods.
@@ IV-A2 Baseline
  • Lightweight Inertial Odometry: IMUNet[11], EfficientNet[11], MobileNet[11], and MnasNet[11].
@@ IV-A3 Metrics
We evaluated performance using Absolute Trajectory Error (ATE) and Relative Trajectory Error (RTE).
@@
ATE measures global trajectory accuracy as the root-mean-square distance between aligned estimated and ground-truth positions:
$ATE=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\|\mathbf{p}_{i}^{\text{est}}-\mathbf{p}_{i}^{\text{gt}}\|^{2}}$ (22)
RTE quantifies local drift by comparing estimated and ground-truth displacements within 1-minute time windows:
$RTE=\sqrt{\frac{1}{M}\sum_{k=1}^{M}\|(\mathbf{p}_{k+\Delta t}^{\text{est}}-\mathbf{p}_{k}^{\text{est}})-(\mathbf{p}_{k+\Delta t}^{\text{gt}}-\mathbf{p}_{k}^{\text{gt}})\|^{2}}$ (23)
IV-A4 Training Details
@@
We adopted a sliding-window strategy, segmenting 200 Hz IMU data into 1-second windows. Training used the Adam optimizer with a batch size of 128, a learning rate of $1\text{e-}4$, and up to 100 epochs. All experiments were conducted on an NVIDIA RTX 4090 GPU.
IV-B Evaluation Across Quadruped Robots and Pedestrians
Table I presents the performance comparison between X-IONet and various baseline methods on the self-collected Go2 quadruped robot dataset. The experimental results show that X-IONet achieved the lowest errors in five out of six test scenarios and obtained the second-lowest error in the Hard 01 scenario, highlighting its robustness to diverse motion patterns of quadruped robots. Furthermore, in terms of average errors across all motion scenarios, X-IONet reduced ATE and RTE by 52.8% and 41.3%, respectively, compared with the strongest baseline method, further confirming its advantage in dynamic state estimation tasks.
@@ IV-B Evaluation Across Quadruped Robots and Pedestrians
Table II presents the performance comparison on the RoNIN pedestrian dataset. The experimental results show that X-IONet achieved state-of-the-art average performance under both Seen and Unseen testing conditions. Compared with the strongest baseline method, X-IONet reduced ATE and RTE by 14.3% and 11.4%, respectively. These results further confirm that the cross-platform design of X-IONet does not compromise performance on pedestrian scenarios, avoiding overfitting to specific gait patterns while maintaining competitive accuracy on human inertial navigation.
Fig. 4 illustrates the trajectory comparisons between X-IONet and baseline methods in several test scenarios. As shown in the figure, X-IONet produces trajectories that are consistently closer to the ground truth in both quadruped robot and pedestrian tasks, demonstrating its superior accuracy in inertial odometry and the effectiveness of its cross-platform design.
IV-C Ablation Study
@@
We conduct ablation experiments to quantify the contribution of each key component in X-IONet.
@@
TABLE III: Results comparison with and without critical components.
Model RoNIN_Seen
ATE RTE
Ours 3.10 2.42
Ours w/o Dual-Stage Attention Mechanism 4.65 ($\uparrow$1.55) 3.37 ($\uparrow$0.95)
Ours w/o Hierarchical Encoder-Decoder Architecture 3.32 ($\uparrow$0.22) 2.63 ($\uparrow$0.21)
Ours w/o Extended Kalman Filter Module 3.41 ($\uparrow$0.31) 2.89 ($\uparrow$0.47)
Ours w/o Uncertainty in Loss Function 3.73 ($\uparrow$0.63) 3.01 ($\uparrow$0.59)
The experimental results in Table 3 show that after removing the Dual-Stage Attention Mechanism, the ATE and RTE metrics increased by 50.0% and 39.3%, respectively. This result confirms that capturing cross-dimensional dependencies via the Dual-Stage Attention is significantly effective in improving the performance of inertial positioning. If the cross-dimensional dependencies are not fully utilized, the accuracy and stability of trajectory prediction will be significantly constrained.
After removing the hierarchical encoder-decoder architecture and applying cross-attention only once in the decoder, the ATE and RTE metrics increased by 7.1% and 8.0%, respectively. This result indicates that the architecture can effectively leverage multi-scale information, thereby enhancing the performance of trajectory prediction.
After removing the Extended Kalman Filter (EKF) module, the ATE and RTE metrics increased by 9.1% and 19.4%, respectively. This result demonstrates that the EKF can effectively optimize the state estimation process, thereby enhancing the robustness of the entire inertial odometry system.
When MSE is used as the loss function instead of the Huber-Gaussian loss, the ATE and RTE metrics increased by 20.3% and 24.4%, respectively. This result indicates that if the displacement prediction network only tries to make the predicted values as close as possible to the true values while ignoring uncertainty modeling during training, the prediction accuracy of the entire framework will be significantly reduced.
IV-D Why the Huber–Gaussian loss?
@@
Table IV compares uncertainty consistency under both loss functions. The TLIO loss is more sensitive to outliers (e.g., sudden IMU disturbances), causing more samples to exceed the $3\sigma$ bound. In contrast, the Huber term attenuates outlier effects, improving coverage by keeping more samples within the $3\sigma$ range. This demonstrates that the Huber–Gaussian loss yields more stable and reliable uncertainty estimates.
IV-E Effectiveness of the Rule-Based Expert Selection Network
Figure 5: The trajectory of the quadruped robot predicted using the pedestrian model.
TABLE V: Comparison of results with and without classifer.
Model RoNIN_Seen ATE RTE Ours 3.10 2.42 Ours w/o Motion Type Classifer 3.47 ($\uparrow$0.37) 2.72 ($\uparrow$0.3)
Human and quadruped motion patterns differ substantially in dynamics and inertial characteristics, causing models trained on one domain to fail on the other. As shown in Fig. 5, applying a human-trained model to quadruped data results in severe trajectory distortion.
We evaluate whether direct mixed-domain training can replace expert selection. After jointly training on pedestrian and quadruped data, performance on pedestrian tests still drops—ATE and RTE degrade by 11.9% and 12.4%, respectively (Table V). These results indicate that naive joint training cannot achieve true cross-platform generalization. In contrast, the proposed rule-based classifier effectively routes IMU data to specialized experts, providing a simple and reliable mechanism for cross-platform inertial odometry.
@@ V CONCLUSIONS
This paper presented X-IONet, a cross-platform inertial odometry framework. By extracting frequency-domain features via discrete wavelet transform and fusing them with temporal signals, the network captures temporal, frequency, inter-variable, and local dependencies. The model regresses displacement and uncertainty, which are refined with an Extended Kalman Filter (EKF) for robust state estimation. Extensive experiments on both the RoNIN pedestrian dataset and a newly collected Unitree Go2 quadruped dataset demonstrated state-of-the-art performance. X-IONet reduced ATE and RTE by 14.3% and 11.4% on RoNIN, and by 52.8% and 41.3% on Go2, compared with the best existing methods. Ablation studies confirmed the effectiveness of the dual-stage attention mechanism, frequency-domain input, and EKF refinement. In addition, a rule based front-end classification module enabled robust cross-platform generalization across humans and robots. Future work will extend this framework to a wider range of robotic platforms and investigate multimodal sensor fusion to further enhance generalization, robustness, and long-term autonomy.

### v2

@@ III-C2 State Propagation
$\hat{b}_{g_{(n+1)}}=\hat{b}_{g_{n}}+n_{g_{n}}$ (13)
$\hat{b}_{a_{(n+1)}}=\hat{b}_{a_{n}}+n_{a_{n}}$ (14)
where $\exp_{SO3}$ is the exponential map for the rotation group, and $\Delta t$ is the sampling period of the inertial data. $\boldsymbol{g}$ represents the gravity acceleration in the world frame. $n$ donates Gaussian-distributed noise.
@@ III-C3 State Update
$h(\boldsymbol{X})=\mathbf{R}^{T}(\mathbf{{}^{w}}\boldsymbol{p}_{j}-\mathbf{{}^{w}}\boldsymbol{p}_{i})=\boldsymbol{\hat{d}}_{ij}+n_{ij}$ (15)
where $\mathbf{R}^{T}$ represents the transformation that converts the displacement difference in the world frame to the local frame. Using the measurement Jacobian matrix $\mathbf{H}$ and the covariance $\hat{\boldsymbol{\Sigma}}_{ij}$ output by the neural network, the Kalman gain $\mathbf{K}$ is calculated as:
@@
where $\oplus$ denotes the addition operation on the $SO(3)$ group when updating pose state variables, while other state variables use ordinary addition. The covariance is updated following the standard EKF correction step.
@@ IV EXPERIMENT
Figure 4: Trajectory comparisons of representative experimental results. The top row corresponds to GrandTour test sequences, and the bottom row corresponds to RoNIN test sequences.
@@ IV-A1 Datasets
We evaluated the proposed method on both pedestrian and quadruped robot inertial datasets.
For pedestrians, we use the RoNIN[15] dataset, a large-scale inertial odometry benchmark collected with handheld smartphones. The dataset contains 42.7 hours of IMU measurements with corresponding ground-truth trajectories. We follow the official training, validation, and test splits. The test set is further divided into Seen and Unseen subsets, each comprising 32 trajectories. The Seen subset includes trajectories from subjects present in the training set, whereas the Unseen subset contains trajectories from previously unseen subjects.
For quadruped robots, we use the public GrandTour dataset[13], which provides synchronized multimodal measurements collected from quadruped platforms in real-world environments. In this work, we use the IMU data from the NovAtel CPT7 device as the inertial input and the corresponding odometry trajectories as ground truth. After data screening, 36 trajectories with complete and reliable IMU and ground-truth data are retained, while the remaining sequences are excluded due to missing ground-truth trajectories. The selected subset is split into 24 training, 6 validation, and 6 test sequences.
In addition to the public dataset, we collected a quadruped inertial dataset using a Unitree Go2 robot. As shown in Fig. 5, an iPhone 14 Pro Max was rigidly mounted on the robot to record IMU data, while the reference trajectory was obtained from onboard LiDAR odometry. The phone’s IMU provided measurements at 100 Hz, which were resampled to 200 Hz along with the ground-truth trajectories using linear interpolation. The dataset consists of 30 sequences with a total duration of approximately 1.5 hours and is used for additional validation under a different platform and sensor configuration. Each sequence lasts approximately 3 minutes, with traveled distances ranging from 50 to 100 m. During steady forward locomotion, the nominal speed is approximately 0.46 m/s.
TABLE I: Results comparison of different models on RoNIN, GrandTour and Go2 dataset.
Model RoNIN_seen RoNIN_unseen GrandTour Go2
ATE (m) RTE (m) ATE (m) RTE (m) ATE (m) RTE (m) ATE (m) RTE (m)
RoNIN-ResNet 4.11 3.28 5.26 4.62 6.51 5.32 3.62 1.95
RoNIN-TCN 4.31 5.02 6.74 5.10 8.21 7.39 3,61 2.46
RONIN-LSTM 6.19 4.56 7.61 5.87 8.56 7.23 3.98 2.08
TLIO 4.74 3.96 6.11 4.75 6.09 4.63 2.22 1.37
IMUNet 6.56 5.90 8.35 6.02 8.03 7.32 3.17 1.77
MobileNet 6.89 5.09 8.80 6.96 7.53 6.48 3.56 2.07
MnasNet 3.96 2.80 6.08 5.03 6.77 6.04 3.45 1.93
EfficientNet 4.59 3.46 6.25 4.84 6.89 5.68 2.18 1.43
CTIN 4.34 3.08 5.25 4.54 6.67 5.65 3.54 2.03
Ours 3.10 2.42 4.79 4.05 5.37 4.18 1.03 0.84
@@ IV-A2 Baseline
  • Lightweight Inertial Odometry: IMUNet[34], EfficientNet[34], MobileNet[34], MnasNet[34].
@@ IV-A3 Metrics
We evaluated performance using Absolute Trajectory Error (ATE) and Relative Trajectory Error (RTE). ATE measures global trajectory accuracy as the root-mean-square distance between aligned estimated and ground-truth positions. RTE quantifies local drift by comparing estimated and ground-truth displacements within 1-minute time windows.
IV-A4 Training and Inference
@@
For a fair comparison, all learning-based baseline methods were retrained on the training split of the corresponding target dataset before evaluation. We adopted a sliding-window strategy, segmenting 200 Hz IMU data into 1-second windows. Training for all models used the Adam optimizer with a batch size of 128, a learning rate of $1\text{e-}4$, and up to 100 epochs. All experiments were conducted on an NVIDIA RTX 4090 GPU. Compared with one end-to-end generalist model, X-IONet activates only a single expert network for each input window during inference, resulting in efficient deployment.
IV-B Evaluation on Pedestrian Datasets
@@
Table I summarizes the performance of all compared methods on the RoNIN pedestrian dataset. X-IONet achieves state-of-the-art performance under both Seen and Unseen testing conditions. Specifically, compared with the strongest baseline, X-IONet reduces ATE and RTE by 14.3% and 11.4%, respectively. These results confirm that the cross-platform design of X-IONet does not compromise accuracy on pedestrian scenarios, effectively capturing human gait patterns without overfitting to specific motion styles.
Fig. 4 (left panel) presents representative trajectories for the RoNIN test sequences. Only two representative baseline methods are visualized: CTIN[29], which employs a Transformer-based architecture, and TLIO[23], which uses an EKF framework. The trajectories estimated by X-IONet remain consistently closer to the ground truth compared to the baselines, qualitatively supporting the quantitative improvements reported in Table I.
IV-C Evaluation on Quadruped Robot Datasets
Quadruped robot evaluation is performed on two datasets: the public GrandTour dataset and our self-collected Go2 dataset. Table I reports the corresponding results. On the GrandTour dataset, X-IONet achieves the best overall performance, reducing ATE and RTE by 11.8% and 9.7% compared with the strongest baseline. These results indicate that X-IONet can accurately model quadruped inertial dynamics rather than relying on patterns learned from pedestrian motion. For our self-collected Go2 dataset, the performance gains are even more pronounced. X-IONet reduces ATE by 52.8% and RTE by 41.3% relative to the strongest baseline. These improvements demonstrate that the proposed framework is robust to different quadruped platforms and sensor configurations. Fig. 5 shows the quadruped robot used in the experiments, while Fig. 6 presents a representative trajectory comparison on a Go2 test sequence. As in the GrandTour evaluation, CTIN and TLIO are included as reference baselines. The trajectories estimated by X-IONet remain closest to the ground truth, visually confirming the quantitative advantages reported in Table I.
Overall, these evaluations demonstrate that X-IONet consistently generalizes across pedestrian and quadruped scenarios, including both public benchmarks and a distinct real-world robot platform. The method effectively captures platform-specific motion dynamics while maintaining robustness across different sensor setups and locomotion modalities.
Figure 5: The quadruped robot used in the experiments.
Figure 6: Trajectory comparison on a representative self-collected Go2 test sequence (left), and the trajectory of the quadruped robot predicted using the pedestrian model (right).
IV-D Ablation Study
@@
We conducted ablation experiments to quantify the contribution of each key component in X-IONet. The following variants were evaluated:
@@
The results in Table II show the quantitative impact of each component across the RoNIN, GrandTour, and Go2 datasets. Removing the Dual-Stage Attention Mechanism consistently degrades performance, with ATE and RTE increasing by 33.4%–107.8% and 34.6%–85.7%, respectively. This highlights the importance of jointly modeling temporal dependencies and cross-dimensional correlations for accurate inertial odometry across pedestrian and quadruped scenarios.
Ablating the hierarchical encoder-decoder architecture leads to ATE and RTE increases of 7.6%–40.8% and 8.6%–20.2%, demonstrating that multi-scale feature aggregation improves the network’s ability to capture both fine-grained motion details and long-range temporal context.
Removing the EKF refinement results in ATE and RTE increases of 12.1%–32.0% and 11.2%–15.7%, confirming that EKF-based integration reduces drift and enhances trajectory accuracy over time. Finally, removing uncertainty-awareness from the loss function increases ATE and RTE by 16.7%–69.9% and 24.4%–61.9%, respectively, indicating that modeling prediction uncertainty significantly improves robustness across diverse motion patterns and sensing conditions.
TABLE II: Ablation study on RoNIN, GrandTour and Go2 datasets.
Model Variation RoNIN GrandTour Go2
ATE (m) RTE (m) ATE (m) RTE (m) ATE (m) RTE (m)
Ours 3.95 3.24 5.37 4.18 1.03 0.84
w/o Dual-Stage Attention 5.27 4.36 7.61 6.32 2.14 1.56
w/o Hierarchical Encoder-Decoder 4.37 3.52 5.78 4.87 1.45 1.01
w/o Extended Kalman Filter 4.46 3.75 6.02 4.65 1.36 0.97
w/o Loss Uncertainty 4.61 4.03 7.19 5.87 1.75 1.36
IV-E Why the Huber–Gaussian loss?
@@
Huber–Gaussian Loss 99.2%
Loss Function in TLIO[23] 98.7%
Table III compares uncertainty consistency under both loss functions. The TLIO loss is more sensitive to outliers (e.g., sudden IMU disturbances), causing more samples to exceed the $3\sigma$ bound. In contrast, the Huber term attenuates outlier effects, improving coverage by keeping more samples within the $3\sigma$ range. This demonstrates that the Huber–Gaussian loss yields more stable and reliable uncertainty estimates.
To further justify the choice of the Huber–Gaussian loss parameters, we evaluated a grid of Huber thresholds and Gaussian regularization weights, with $\delta\in\{0.001,0.005,0.01\}$ and $\lambda\in\{0.0001,0.001,0.01\}$. As shown in Fig. 7, we measured the RMSE of the predicted displacement and the $3\sigma$ coverage. Among all tested settings, $\delta=0.005$ and $\lambda=0.0001$ achieved the best overall performance.
Figure 7: Parameter sensitivity analysis of the Huber-Gaussian loss on the GrandTour dataset.
IV-F Effectiveness of the Rule-Based Expert Selection Network
TABLE IV: Classification accuracy of the rule-based expert selection network evaluated on 1-second windows from the RoNIN pedestrian dataset and the GrandTour quadruped robot dataset.
Dataset Target Class Number of Samples Accuracy
RoNIN Pedestrian 376,262 99.72%
GrandTour Quadruped Robot 40,445 100.00%
Human and quadruped motion patterns differ substantially in dynamics and inertial characteristics, causing models trained on one domain to fail on the other. As shown in Fig. 6, applying a human-trained model to quadruped data results in severe trajectory distortion. We tested the classifier on 376,262 one-second windows from the pedestrian RoNIN dataset and 40,445 windows from the quadruped robot GrandTour dataset. As summarized in Table IV, the classifier achieved accuracies of 99.72% and 100.00%, respectively, demonstrating highly reliable platform recognition and expert routing performance.
Figure 8: Boxplot comparison of ATE and RTE for X-IONet with and without the motion type classifier. The mean values are annotated above each boxplot.
We evaluate whether direct mixed-domain training can replace expert selection. After jointly training on pedestrian and quadruped data, performance on pedestrian tests still drops—ATE and RTE degrade by 11.9% and 12.4%, respectively (Fig. 8). These results indicate that naive joint training cannot achieve true cross-platform generalization. In contrast, the proposed rule-based classifier effectively routes IMU data to specialized experts, providing a simple and reliable mechanism for cross-platform inertial odometry.
@@ V CONCLUSIONS
This paper presented X-IONet, a cross-platform inertial odometry framework. The displacement prediction network employs a dual-stage attention architecture to jointly capture long-range temporal dependencies and inter-axis correlations from raw inertial signals. The model regresses displacement and uncertainty, which are refined with an EKF for robust state estimation. Extensive experiments on both the RONIN pedestrian dataset and a newly collected Unitree Go2 quadruped dataset demonstrated state-of-the-art performance. X-IONet reduced ATE and RTE by 14.3% and 11.4% on RoNIN, and by 52.8% and 41.3% on Go2, compared with the best existing methods. Ablation studies confirmed the effectiveness of the dual-stage attention mechanism, the hierarchical encoder-decoder architecture, and the EKF refinement. In addition, a rule-based front-end classification module enabled robust cross-platform generalization across humans and robots.
The current evaluation is limited to trained pedestrian and quadruped domains; performance on other platforms, unseen robots, and heterogeneous IMU configurations remains to be explored. Future work will extend the framework to a broader range of platforms and improve robustness to unseen carrier instances, different robot models, and diverse IMU hardware. An important direction is to investigate whether cross-robot generalization is better achieved through jointly trained experts or through robot-specific experts combined with a more fine-grained classifier.